OpenStudy (anonymous):

Solve the polynomial inequality. Express the solution set in interval notation. 25x^2+1is greater than or equal to -10x

OpenStudy (hoblos):

\[25x ^{2} +10x +1 = (5x+1)^{2} \ge 0 \forall x\]

OpenStudy (anonymous):

i dont understand

OpenStudy (hoblos):

\[25x ^{2} +1 \ge -10x\] =>\[25x ^{2} +10x +1 \ge 0\] =>\[(5x+1)^{2} \ge 0\] and this is always true for any x

OpenStudy (hoblos):

get it ?

OpenStudy (anonymous):

yes thank you

OpenStudy (hoblos):

:D